### v1

# Absolutely Maximal Contextual Correlations

Abstract
The foundational work by Bell led to an interest in understanding non-local correlations that arise from entangled states shared between distinct, spacelike-separated parties, which formed a foundation for the theory of quantum information processing. We investigate the question of maximal correlations analogous to the maximally entangled states defined in the entanglement theory of multipartite systems. To formalize this, we employ the sheaf-theoretic framework for contextuality, which generalizes non-locality. This provides a new metric for correlations called contextual fraction (CF), which ranges from 0 (non-contextual) to 1 (maximally contextual). Using this, we have defined the absolutely maximal contextual correlations (AMCC), which are maximally contextual and have maximal marginals, which captures the notion of absolutely maximal entangled (AME) states. The Popescu-Rohrlich (PR) box serves as the bipartite example, and we construct various extensions of such correlations in the tripartite case. An infinite family of various forms of AMCC is constructed using the parity check method and the constraint satisfiability problem (CSP) scheme. We also demonstrate the existence of maximally contextual correlations, which do not exhibit maximal marginals, and refer to them as non-AMCC. The results are further applied to secret sharing and randomness extraction using AMCC correlations.
I Introduction
The remarkable analysis by Einstein, Podolsky, and Rosen (EPR) Einstein et al. (1935) revealed the peculiar behavior of nature and questioned the completeness of quantum mechanics. The quantum mechanical state used in their argument exhibits nonclassical correlations, which Schrodinger later termed entanglement. Attempts to address the questions raised by EPR culminated in the seminal work of Bell Bell (1964), who developed a theoretical framework to test the foundations of quantum theory. Bell’s theorem demonstrates the incompatibility of quantum mechanical predictions with any theory that combines the assumptions of determinism, locality, and measurement independence.
The exploration of entanglement has led to profound insights and practical advantages, underpinning the entire field of quantum information theory and quantum computation. A foundational framework for this study treats entanglement as a resource that can be manipulated through local operations and classical communication (LOCC) Horodecki et al. (2009). For bipartite systems, asymptotic entanglement transformation via LOCC has been extensively studied, along with measures of pure-state bipartite entanglement Popescu and Rohrlich (1997); Nielsen (1999). From these works, the concept of a maximally entangled state in bipartite systems is well established Nielsen (1999); Popescu and Rohrlich (1997); Dür et al. (2000). The structure becomes richer for multipartite systems. Even for three qubits, two inequivalent classes exist under stochastic LOCC (SLOCC) Dür et al. (2000), requiring new measures to quantify genuine multipartite entanglement. In general, the concepts of entanglement transformation, quantification, and maximal entanglement become intractable for multipartite systems due to the exponential complexity in the number of parameters. Consequently, despite extensive research Verstraete et al. (2002); Kraus (2010); de Vicente et al. (2013); Miyake (2003); Gour and Wallach (2011); Spee et al. (2017); Li et al. (2024); Gunn et al. (2023); Ma et al. (2011); Xie and Eberly (2021), fundamental questions remain open: Does a unambiguous concept of a “maximally entangled state” exist for multipartite systems, and if so, how should it be defined?
The correlations obtained by measuring an entangled state constitute another important form of nonclassicality, both from a resource-theoretic and an applied perspective. In a Bell experiment, these correlations can be viewed abstractly as a “black box” characterized by inputs and outputs, with its internal physical structure being irrelevant. This abstraction makes them a powerful resource for device-independent protocols Pironio et al. (2010); Moreno et al. (2020). Correlations that violate a Bell inequality are termed nonlocal and belong to a strictly larger set than those admitting a local hidden-variable model. Tsirelson established that the quantum maximum for the paradigmatic Clauser-Horne-Shimony-Holt (CHSH) expression is $2\sqrt{2}$ Cirel’son (1980). More generally, post-quantum nonlocal correlations—supersets of the quantum set—have also been studied, as exemplified by the Popescu-Rohrlich box Popescu and Rohrlich (1994).
The theories of entanglement and nonlocal correlations are deeply intertwined, presenting significant conceptual intricacies Schmid et al. (2023). This interplay is clarified by the framework of local operations and shared randomness (LOSR) Schmid et al. (2020); Rosset et al. (2020), which serves as the correlational analog to LOCC in entanglement theory. To understand genuine multipartite correlations, a more general operational approach employs wiring and classical communication prior to inputs (WCCPI) Gallego et al. (2012), demonstrating that earlier definitions were incomplete Bancal et al. (2011). A fundamental difficulty arises from a key divergence: while the classification of pure entangled states into equivalence classes under LOCC and SLOCC is well-established, no analogous structure exists for nonlocal correlations. In the bipartite case, LOCC yields a single equivalence class of maximally entangled states—a definitive “golden unit.” However, no corresponding unique unit has been identified for correlations. Concepts such as single-copy or asymptotic distillation do not translate directly from entanglement theory, preventing a straightforward mapping from maximal entanglement to maximal correlation. Consequently, many correlation sets are incomparable in the asymptotic limit Ghosh et al. (2024). Thus, a central challenge is to formulate rigorous definitions for maximal entanglement and maximal correlations in genuine multipartite systems, an area that remains actively unresolved.
Correlations from an experimental setup can be viewed more generally using a sheaf-theoretic approach Abramsky and Brandenburger (2011). Any dataset described by $(n,m,o)$, where $n$ is the number of parties, $m$ is the number of measurements per party, and $o$ is the set of possible outcomes, can be formalized as a measurement scenario. In this framework, compatible sets of measurements (called contexts) and their possible outcomes are organized so that probability distributions over outcomes define the empirical model.
A no-signaling empirical model is one in which outcome probabilities do not depend on which other measurements are performed, generalized by the compatibility condition. If these probabilities can be derived from a single global assignment, the model is noncontextual; otherwise, it is contextual. More formally, any empirical model can be represented as a linear system: if this system admits a non-negative solution, the model is noncontextual; otherwise, it is contextual. In the possibilistic case, a stronger criterion defines the notion of a strongly contextual model. Any empirical model can be expressed as a convex combination of a noncontextual model (with coefficient non-contextual fraction, $\mathrm{NCF}=p$) and a strongly contextual model (with coefficient contextual fraction $\mathrm{CF}=1-p$). If CF=1 ($p=0$), the model is maximally contextual; if CF=0 ($p=1$), the model is noncontextual Abramsky et al. (2017). It has been shown that a model is strongly contextual if and only if it is maximally contextual Abramsky and Brandenburger (2011).
These results and framework has broader implications for our understanding of computational advantage. In particular, within quantum computation, this advantage manifests as the ability to transition from limited computational power, characterized by the parity complexity class, to full quantum universality Anders and Browne (2009). This transition has been demonstrated using GHZ-type correlations and Mermin-type arguments. Moreover, the contextuality has been interpreted in various ways Raussendorf (2013); Okay et al. (2017); Abramsky et al. (2012); Raussendorf (2019); Frembs et al. (2018); Walleghem et al. (2024); Raussendorf et al. (2023); Okay and Raussendorf (2020); Xu et al. (2020); Saha et al. (2019); Beer and Osborne (2018).
It is evident that contextuality is a more general phenomenon than nonlocality, and that correlations exhibiting maximal contextuality are termed maximal contextual correlations, as they do not contain any local fraction. To extend the idea of genuine multipartite entanglement, a special class of entanglement is defined as absolutely maximally entangled (AME) states Helwig et al. (2012). These pure states are maximally entangled, such that for any bipartite reduction, they yield maximally mixed states. In this back drop, we would like to ask the question what is the analogues of AME in the theory of correlations. In this work, we extend this to genuine multipartite correlations and define a new class of correlations as absolutely maximal contextual correlations (AMCC). These correlations represent the maximal contextual correlations, which are also maximally random marginals. In the $(2,2,2)$ scenario, there are exactly eight such AMCCs which are Popescu-Rohrich (PR) boxes. However, in the multipartite scenario like $(3,2,2)$, there are infinitely many such correlations that qualify as AMCCs and reside on the face of the polytope. The mathematical construction of these correlations can be accomplished using the parity check method or the constraint satisfiability problem (CSP) approach. Apart from AMCCs, there exists another class of correlations that are maximally contextual but not maximally random marginals; these are constructed exclusively using the CSP method. The AMCCs correspond to correlations that are analogous to absolutely maximally entangled (AME) states. This property is of major significance in device-independent randomness extraction due to the maximal marginal property, and such correlations can also be utilized in secure communication protocols.
The paper is organized into the following sections. In Section(II), we will discuss the mathematical formalism of correlations and the corresponding convex polytope structure. It also incorporates the mathematical tool of sheaf theory and several useful definitions to facilitate an understanding of contextuality and its various forms. In Section(III), we define the AMCC and present some evidence based on it. Later in the section, we provide the construction methods. We propose two methods here, based on the structure of correlations. We employed two methods: the all-vs-nothing proof and the constraint satisfiability problem (CSP). Lastly, in section (V), we have shown some remarkable applications, demonstrating how the AMCC will be used for randomness extraction and secret sharing purposes.
II Formalism
We begin with a simple picture: two observers, Alice and Bob, share either a classical or quantum state, and each can perform local measurements on their own systems. Alice can choose between two dichotomous measurements, $\{X_{1},X_{1}^{\prime}\}$, and Bob likewise chooses from $\{X_{2},X_{2}^{\prime}\}$. Each measurement yields one of two possible outcomes, labeled 0 or $1$. The experimental statistics are fully described by the set of joint probabilities $p(x_{1},x_{2}\mid X_{1},X_{2})$, giving the probability that Alice obtains outcome $x_{1}$ and Bob obtains outcome $x_{2}$ when they choose measurements $X_{1}$ and $X_{2}$, respectively. This setup is often described as a black-box scenario: each observer holds a device to which they input a measurement choice ($X_{1}$ or $X_{1}^{\prime}$ for Alice, $X_{2}$ or $X_{2}^{\prime}$ for Bob) and receives an outcome ($x_{1}$ or $x_{2}$) in return. The internal workings of the devices are unknown; only the input-output statistics are accessible. This is known as the $(2,2,2)$ Bell scenario, where the triple denotes two parties, each with two measurement choices and two outcomes per measurement.
Consider an $(n,m,o)$ multipartite Bell scenario, where $n$ denotes the number of parties indexed by $i\in\{1,\dots,n\}$, for each party $i$ there are $m_{i}$ measurement settings denoted $X_{i}\in\{1,\dots,m_{i}\}$, and for each measurement $X_{i}=j$ of party $i$ there are $o_{ij}$ possible outcomes denoted $x_{i}\in\{1,\dots,o_{ij}\}$. The statistics are described by the joint conditional probability distribution $p(x_{1},\dots,x_{n}\mid X_{1},\dots,X_{n})$, which must satisfy: normalization, $\sum_{x_{1},\dots,x_{n}}p(x_{1},\dots,x_{n}\mid X_{1},\dots,X_{n})=1$ for all $X_{1},\dots,X_{n}$; positivity, $p(x_{1},\dots,x_{n}\mid X_{1},\dots,X_{n})\geq 0$ for all outcomes and settings; and the no-signaling condition, which requires that for each party $j$ the marginal distribution of the remaining parties is independent of party $j$’s measurement choice:
$\sum_{x_{j}}p(x_{1},\cdots,x_{j},\cdots,x_{n}\mid X_{1},\cdots,X_{j},\cdots,X_{n})=\sum_{x_{j}}p(x_{1},\cdots,x_{j},\cdots,x_{n}\mid X_{1},\cdots,X_{j}^{\prime},\cdots,X_{n})$ (1)
for all $j$, all outcomes $x_{1},\dots,x_{j-1},x_{j+1},\dots,x_{n}$, and all settings $X_{1},\dots,X_{j-1},X_{j},X_{j}^{\prime},X_{j+1},\dots,X_{n}$. The set of probability distributions satisfying these constraints forms a convex polytope known as the no-signaling polytope, with dimension
$\dim P=\prod_{i=1}^{n}\left(\sum_{j=1}^{m_{i}}(o_{ij}-1)+1\right)-1.$
The correlations are called local if they can be reproduced by a local hidden-variable model, i.e.,
$p(x_{1},\dots,x_{n}|X_{1},\dots,X_{n})=\int d\lambda\,\rho(\lambda)\,p(x_{1}|X_{1},\lambda)\cdots p(x_{n}|X_{n},\lambda),$ (2)
where $\rho(\lambda)\geq 0$, $\int d\lambda\,\rho(\lambda)=1$, and $p(x_{i}|X_{i},\lambda)$ denotes the probability of obtaining outcome $x_{i}$ for measurement $X_{i}$ given the hidden variable $\lambda$. Here $\lambda$ represents a set of pre-assigned values for all possible measurement outcomes. The corresponding deterministic probability vectors are defined as
$p^{\lambda}(x_{1},\dots,x_{n}|X_{1},\dots,X_{n})=\begin{cases}1,&\text{if }\forall i,\;\lambda_{iX_{i}}=x_{i},\;i\in\{1,\dots,n\},\\ 0,&\text{otherwise}.\end{cases}$ (3)
The set of local correlations $P$ is the convex hull of these deterministic points, forming the local polytope
$P=\left\{p\in\mathbb{R}^{r}\;|\;p=\sum_{\lambda\in\Lambda}\rho(\lambda)p^{\lambda},\;\rho(\lambda)\geq 0,\;\sum_{\lambda\in\Lambda}\rho(\lambda)=1\right\}.$
For the Bell polytope $P\subseteq\mathbb{R}^{r}$, consider an inequality $\mathbf{a}\cdot\mathbf{p}\geq a_{0}$, where $(\mathbf{a},a_{0})\in\mathbb{R}^{r+1}$. This inequality is called a valid Bell inequality for $P$ if it holds for all $\mathbf{p}\in P$. To verify validity, it suffices to check that it is satisfied by all extreme points of $P$, i.e., by the deterministic vectors. Given such an inequality, the corresponding face of the polytope is defined by the equality condition $F=\{\mathbf{p}\in P\;|\;\mathbf{a}\cdot\mathbf{p}=a_{0}\}$. A face with dimension exactly one less than that of the full polytope is called a facet Pironio (2005). Thus, for any proper face $F$ (with $F\neq\varnothing$ and $F\neq P$), the dimension satisfies $\dim F\leq\dim P-1$. These proper faces, which also satisfy the no-signaling and normalization conditions, are defined by zero constraints:
$F^{NS}=\{\mathbf{p}\in\mathbf{NS}\;|\;p=0\},$ (4)
where $p$ is an element of $\mathbf{p}=\{p(x_{1},\cdots,x_{i}\mid X_{1},\cdots,X_{i})_{i\leq\dim P-1}\}$, and $\mathbf{NS}\subset\mathbb{R}^{r}$ denotes the set of no-signaling correlations. By imposing these zero constraints, the correlations are brought to their maximally reduced form, corresponding to the lowest-dimensional face. The resulting face correlations yield a set of deterministic constraints that cannot be satisfied simultaneously by any global assignment. In other words, this construction leads directly to AVN (all-versus-nothing) proofs (see the Appendix(A)). Hence, all the zeros define a face of the no-signaling polytope, and every deterministic local vertex $\mathbf{p}^{L}\in P$ violates at least one zero constraint. This implies that $F^{NS}\cap\mathbf{p}^{L}=\phi$; in other words, the correlations on the face $F^{NS}$ in Eq.(4) do not contain any local components Liu et al. (2024). These face correlations are referred to as maximal correlations, which are discussed in the next section.
II.1 Contextual correlations
The Bell-scenario $(n,m,o)$ is an empirical dataset obtained in an experiment. Such an empirical model is described using a measurement scenario, which is a triplet $\langle\mathcal{X},\mathcal{M},\mathcal{O}\rangle$, where $\mathcal{X}$ is a set of all observables, $\mathcal{M}$ is a measurement cover, and $\mathcal{O}$ is a set of all the outcomes obtained from each measurement. We will elaborate on these definitions in the following paragraphs.
Definition 1 (Measurement Cover).
The measurement cover is a set $\mathcal{M}\subseteq\mathcal{X}$ such that:
  1. 1.
    $\bigcup\mathcal{M}=\mathcal{X}$
  2. 2.
    The chain condition: if $\mathcal{C},\mathcal{C}^{\prime}\in\mathcal{M}$ and $\mathcal{C}\subseteq\mathcal{C}^{\prime}$, then $\mathcal{C}=\mathcal{C}^{\prime}$.
For example, let us take the $(2,2,2)$-Bell scenario, as introduced earlier. We can write the measurement set:
$\mathcal{X}=\{X_{1},X_{1}^{\prime},X_{2},X_{2}^{\prime}\},$ (5)
the measurement cover:
$\mathcal{M}=\{\{X_{1},X_{2}\},\{X_{1},X_{2}^{\prime}\},\{X_{1}^{\prime},X_{2}\},\{X_{1}^{\prime},X_{2}^{\prime}\}\},$ (6)
and the outcome set $\mathcal{O}=\{0,1\}$. The joint outcome corresponding to each context $\mathcal{C}\in\mathcal{M}$ is called a section $\mathrm{s}$.
Definition 2 (section).
A section $\mathrm{s}$ is a mapping from $\mathcal{U}\subseteq\mathcal{M}\subseteq\mathcal{X}$ to $\mathcal{O}$, i.e.,
$\mathrm{s}:\mathcal{U}\to\mathcal{O}.$ (7)
The set of all sections is defined as the sheaf of events (or a presheaf) $\mathcal{E}:P(X)^{\text{op}}\to\mathsf{Set}$, which is a functor from the opposite category of the poset $P(X)$ to the category of sets Abramsky and Brandenburger (2011). Considering the previous example, for $\mathcal{C}_{1}=\{X_{1},Y_{1}\}\in\mathcal{M}$, the set of events (sections) is
$\mathcal{E}(\mathcal{C}_{1})=\{(0,0),(0,1),(1,0),(1,1)\}.$ (8)
Weight assignments to the sheaf of events are defined via a functor
$\mathcal{D}_{\mathcal{R}}\mathcal{E}:P(X)\to\mathsf{Set},$ (9)
where $\mathcal{D}_{\mathcal{R}}:\mathsf{Set}\to\mathsf{Set}$ is the distribution functor that sends a set $\mathcal{E}$ to
$\mathcal{D}_{\mathcal{R}}(\mathcal{E})=\{\mathcal{P}:\mathcal{E}\to\mathbb{R}_{\geq 0}\mid\sum_{x\in\mathcal{E}}\mathcal{P}(x)=1\},$
and $\mathcal{R}=\{\mathbb{R}_{\geq 0},+,\cdot,0,1\}$ is a semiring of non‑negative reals. Applying this functor to the sheaf $\mathcal{E}$ yields a presheaf of probability distributions.
If $\mathrm{e}\in\mathcal{D}_{\mathcal{R}}(\mathcal{E}(U^{\prime}))$ is a distribution over sections of $U^{\prime}\subseteq\mathcal{X}$, then there is a restriction map $\text{res}_{U}^{U^{\prime}}$ for distributions such that
$\begin{split}\mathcal{D}_{\mathcal{R}}(\mathcal{E}(U^{\prime}))&\xrightarrow{\text{res}_{U}^{U^{\prime}}}\mathcal{D}_{\mathcal{R}}(\mathcal{E}(U)),\\ \mathrm{e}&\mapsto\mathrm{e}|_{U}.\end{split}$ (10)
Here $\mathrm{e}|_{U}$ is obtained by summing over all $\mathrm{s}^{\prime}\in\mathcal{E}(U^{\prime})$ that restrict to a given $\mathrm{s}\in\mathcal{E}(U)$:
$\mathrm{e}|_{U}(\mathrm{s})=\sum_{\begin{subarray}{c}\mathrm{s}^{\prime}\in\mathcal{E}(U^{\prime})\\ \mathrm{s}^{\prime}|_{U}=\mathrm{s}\end{subarray}}\mathrm{e}(\mathrm{s}^{\prime}).$ (11)
Thus $\mathrm{e}|_{U}$ is the marginal distribution of $\mathrm{e}$ over the subset $U$.
Definition 3 (Empirical model).
Given a measurement scenario $\langle\mathcal{X},\mathcal{M},\mathcal{O}\rangle$, an empirical model is a family of distributions $\{\mathrm{e}_{\mathcal{C}}\}_{\mathcal{C}\in\mathcal{M}}$, where each $\mathrm{e}_{\mathcal{C}}\in\mathcal{D}_{\mathcal{R}}(\mathcal{E}(\mathcal{C}))$.
Definition 4 (Generalized no‑signaling).
Let $\mathcal{C}_{1},\mathcal{C}_{2}\in\mathcal{M}$ with $\mathrm{e}_{\mathcal{C}_{1}}\in\mathcal{D}_{\mathcal{R}}\mathcal{E}(\mathcal{C}_{1})$ and $\mathrm{e}_{\mathcal{C}_{2}}\in\mathcal{D}_{\mathcal{R}}\mathcal{E}(\mathcal{C}_{2})$. The generalized no‑signaling condition is
$\mathrm{e}_{\mathcal{C}_{1}}|_{\mathcal{C}_{1}\cap\mathcal{C}_{2}}=\mathrm{e}_{\mathcal{C}_{2}}|_{\mathcal{C}_{1}\cap\mathcal{C}_{2}}.$ (12)
An empirical model satisfying this compatibility condition is called a no‑signaling empirical model.
The set of all sections over $\mathcal{X}$, i.e. $\mathcal{E}(\mathcal{X})=\mathcal{O}^{\mathcal{X}}$, are called global sections. A global section is a map $g:\mathcal{X}\to\mathcal{O}$, and a distribution over these sections is called a global distribution $\mathrm{e}_{\mathcal{X}}\in\mathcal{D}_{\mathcal{R}}\mathcal{E}(\mathcal{X})$, satisfying
$\mathrm{e}_{\mathcal{C}}(\mathrm{s}_{\mathcal{C}})=\sum_{\begin{subarray}{c}\mathrm{s}^{\prime}\in\mathcal{E}(\mathcal{X})\\ \mathrm{s}^{\prime}|_{\mathcal{C}}=\mathrm{s}_{\mathcal{C}}\end{subarray}}\mathrm{e}_{\mathcal{X}}(\mathrm{s}^{\prime}),$ (13)
i.e., $\mathrm{e}_{\mathcal{C}}$ is the marginal of $\mathrm{e}_{\mathcal{X}}$. Hence, “If such a global probability distribution exists, the model corresponds to a deterministic hidden‑variable scenario.”
The existence of a global section implies that the model can be described by a non‑contextual hidden‑variable theory. One may ask whether, for a given empirical model, a global section exists. This question can be addressed using linear algebra. For a measurement scenario $\langle\mathcal{X},\mathcal{M},\mathcal{O}\rangle$ with $n$ global sections and $m$ local sections, the restriction map can be represented as an incidence matrix $M]_{n\times m}$. Moreover, if a global distribution $\mathbf{d}\in\mathcal{D}_{\mathcal{R}}\mathcal{E}(\mathcal{X})$ exists, then the following linear system must have a solution over the semiring of non‑negative reals:
$\mathbf{M}\,\mathbf{d}=\mathbf{v},\qquad\mathbf{d}\geq\mathbf{0},$ (14)
where $\mathbf{M}$ is the incidence matrix, $\mathbf{d}$ is the global probability distribution, and $\mathbf{v}$ is the vector of local probability assignments. The problem reduces to determining whether a non‑negative solution $\mathbf{d}$ exists.
Definition 5 (Contextual).
A no‑signaling empirical model $\{\mathrm{e}_{\mathcal{C}}\}_{\mathcal{C}\in\mathcal{M}}$ is contextual if and only if the linear system in Eq. (14) admits no solution over the non‑negative reals.
If one replaces the semiring of non‑negative reals by the Boolean semiring, the probabilistic empirical model reduces to a possibilistic model.
Definition 6 (Strong contextuality).
Given a possibilistic model derived from an empirical model, if there exists no global section $g$ such that $\forall\mathcal{C}\in\mathcal{M},\;\mathrm{e}_{\mathcal{C}}(g|_{\mathcal{C}})>0$, then the model is called strongly contextual.
Any empirical model can be decomposed into the following convex combination:
$M=pM^{\mathrm{NC}}+(1-p)M^{\mathrm{SC}},$ (15)
where the noncontextual fraction $\mathrm{NCF}=p$ and the contextual fraction $\mathrm{CF}=1-p$. If $p=0$, the model is maximally contextual; if $p=1$, the model is noncontextual. It has been shown that a model is strongly contextual if and only if it is maximally contextual Abramsky et al. (2017). The contextual fraction can be computed via linear programming (specifically, using the simplex method) Lee (2016). This involves setting up an objective function and linear constraints, which together define a multidimensional convex polytope. The simplex method provides an efficient algorithm for such optimization problems, relying on three main components: slack variables, a tableau (data table), and pivot operations. Thus, it is well-suited for calculating the CF value.
Since non-locality is a special case of contextuality, and strong contextuality represents the maximal form of contextuality, the correlations associated with it are referred to as maximal contextual correlations. The extension of such a correlation to a multipartite scenario has yet to be explored.
The correlations corresponding to maximal contextuality lie on the face of the nonlocal polytope, as they do not contain any local fractions  [^1] Any empirical model can be decomposed as a convex mixture of contextual and noncontextual models, with contextuality being a more general concept than nonlocality. The convex decomposition of a nonlocal correlation can be expressed as $p(x_{1},\ldots,x_{n})=q^{L}p_{L}(x_{1},\ldots,x_{n})+q^{NL}p_{NL}(x_{1},\ldots,x_{n}),$ where (pL) denotes the local part and (pNL) the nonlocal part. This decomposition is a special case within the broader framework of contextuality. Hence, the condition that the contextual fraction $CF=1$ corresponds to a special case where the nonlocal fraction is also maximal. This is significant because such correlations also correspond to all-versus-nothing (AVN) proofs, such as Mermin’s arguments Liu et al. (2024). In the probabilistic form of AVN, the presence of zeros determines the structure of these correlations; these zeros specify the facets that maximally reduce the dimension of the face while still satisfying the no-signaling and normalization conditions. We utilize this property in Section III for construction purposes. While the concept of maximal entanglement is well established for quantum states, its direct analogue in terms of correlations remains somewhat ambiguous. To clarify the notion of maximality in correlations, we take absolutely maximally entangled (AME) states Helwig et al. (2012) as a reference point, since their genuine bipartite marginals are maximally mixed. Accordingly, we focus on correlations that are not only maximally contextual but also exhibit maximal marginality. This approach leads to the identification of a new class of correlations and provides a precise definition of maximality in the context of correlations, which will be introduced in the next section.
III Absolutely maximal contextual correlations
The results and calculations presented here are obtained using the mathematical framework of sheaf theory Abramsky and Brandenburger (2011), which has been employed to define AMCCs. Contextuality—a fundamentally nonclassical property—serves as a key feature within this model and is quantified by the contextual fraction ($CF$), which ranges between 0 and $1$. Accordingly, three classes of models can be distinguished: noncontextual $(\mathrm{CF}=0)$, contextual $(0<\mathrm{CF}<1)$, and maximally contextual $(\mathrm{CF}=1)$ Abramsky et al. (2017). Any empirical model can thus be expressed as a convex combination of a noncontextual and a maximally contextual component.
Furthermore, the mathematical formalism allows for the simulation of empirical models and their associated measurement scenarios, as defined in Section II. In this section, we utilize these concepts to introduce a new class of maximal contextual correlations, based on the maximally contextual model. These correlations are of particular interest as they satisfy the all-versus-nothing (AVN) proof, and are therefore referred to as face-non-signaling (FNS) correlations Liu et al. (2024). The required mathematical definitions are defined below in this Section.
Definition 7 (Maximal Marginals).
For any $(n,m,o)$ Bell scenario, the correlations are called maximal marginals if all marginals satisfy
$e|_{U}=p(x_{1},\dots,x_{k}\mid X_{1},\dots,X_{k})=\frac{1}{2^{k}}\quad\forall\;k<n,$ (16)
where $U=\{X_{1},\dots,X_{k}\}\subset\mathcal{X}$ and each $x_{i}\in\{0,1\}$.
Definition 8 (Absolutely Maximally Contextual Correlations).
An $(n,m,o)$ correlation is called absolutely maximally contextual if it is simultaneously maximally contextual and maximal marginal.
The bipartite case of AMCCs is represented by the PR-box Popescu and Rohrlich (1994) for the $(2,2,2)$-Bell scenario, for which the contextual fraction is $\mathrm{CF}=1$. The Bell polytope $P$ for this scenario has dimension $8$ and contains $24$ vertices: $16$ local vertices and $8$ PR-box vertices. The eight PR boxes are given by Barrett et al. (2005)
$P_{\text{PR}}^{\alpha\beta\gamma}(x_{1}x_{2}|X_{1}X_{2})=\begin{cases}\frac{1}{2},&\begin{aligned} x_{1}\oplus x_{2}&=X_{1}X_{2}\oplus\alpha X_{1}\\ &\oplus\beta X_{2}\oplus\gamma,\end{aligned}\\ 0,&\text{otherwise},\end{cases}$ (17)
where $\alpha,\beta,\gamma\in\{0,1\}$ and $\oplus$ denotes addition modulo $2$. Defining $f=X_{1}X_{2}\oplus\alpha X_{1}\oplus\beta X_{2}\oplus\gamma$, the PR box can also be expressed as
$P_{\text{PR}}^{\alpha\beta\gamma}(x_{1}x_{2}|X_{1}X_{2})=\frac{1}{2}\left(\delta_{0}^{x_{1}}\delta_{f}^{x_{2}}+\delta_{1}^{x_{1}}\delta_{f\oplus 1}^{x_{2}}\right),$ (18)
where $\delta_{a}^{b}$ is the Kronecker delta. The bipartite marginals for the correlation are maximal marginals:
$\begin{split}&P(x_{1}|X_{1})=\frac{1}{2}\left(\delta_{0}^{x_{1}}+\delta_{1}^{x_{1}}\right),\\ &P(x_{2}|X_{2})=\frac{1}{2}\left(\delta_{0}^{x_{2}}+\delta_{1}^{x_{2}}\right)\end{split}$ (19)
Since the marginals corresponding to these PR box correlations are maximal marginals, and the contextual fraction is $\mathrm{CF}=1$, for the $(2,2,2)$-Bell scenario, the PR boxes are the AMCCs.
III.1 $(3,2,2)$-Bell scenario
Consider the three-party $(3,2,2)$-Bell scenario. The corresponding no-signaling polytope has dimension $26$ with $46$ extremal points, which can be grouped into three categories: local, two-way nonlocal, and three-way nonlocal Barrett et al. (2005). We investigate which types of correlations in $(3,2,2)$ Bell scenarios give rise to Absolutely Maximally Contextual Correlations (AMCCs).
In the $(3,2,2)$ scenario, there are $8$ measurement contexts, and each context admits $8$ possible outcome assignments (sections), yielding a total of $64$ local joint assignments. Over the full set of observables $\mathcal{X}=\{X_{1},X_{1}^{\prime},X_{2},X_{2}^{\prime},X_{3},X_{3}^{\prime}\}$, there are $2^{6}=64$ possible global assignments. Whether such global sections exist can be checked using the linear programming approach described earlier.
As an example, consider the correlation obtained from the GHZ state
$|\text{GHZ}\rangle=\frac{|000\rangle+|111\rangle}{\sqrt{2}}$
with each party performing measurements in the $X$ or $Y$ basis (e.g., $X_{1}=\sigma_{x}$, $X_{1}^{\prime}=\sigma_{y}$, and similarly for parties 2 and 3). The resulting probability distribution is given by
$P(x_{1}x_{2}x_{3}|X_{1}X_{2}X_{3})=\begin{cases}\frac{1}{4},&x_{1}\oplus x_{2}\oplus x_{3}=1\oplus X_{1}X_{2}X_{3}\oplus X_{1}X_{2}\oplus X_{2}X_{3}\oplus X_{3}X_{1},\\[6.0pt] \frac{1}{8},&\text{for }(X_{1},X_{2},X_{3}^{\prime}),\;(X_{1},X_{2}^{\prime},X_{3}),\;(X_{1}^{\prime},X_{2},X_{3}),\;(X_{1}^{\prime},X_{2}^{\prime},X_{3}^{\prime}),\\[6.0pt] 0,&\text{otherwise},\end{cases}$ (20)
where $X_{i},X_{i}^{\prime}\in\{0,1\}$ encodes the measurement choices (e.g., 0 for $X$, 11 for $Y$) and $x_{i}\in\{0,1\}$ are the outcomes. The bipartite marginals of this correlation are uniformly random:
$\begin{split}p(x_{1}x_{2}|X_{1}X_{2})&=\frac{1}{4}\left[\delta_{0}^{x_{1}}\delta_{0}^{x_{2}}+\delta_{0}^{x_{1}}\delta_{1}^{x_{2}}+\delta_{1}^{x_{1}}\delta_{0}^{x_{2}}+\delta_{1}^{x_{1}}\delta_{1}^{x_{2}}\right],\\ p(x_{1}x_{3}|X_{1}X_{3})&=\frac{1}{4}\left[\delta_{0}^{x_{1}}\delta_{0}^{x_{3}}+\delta_{0}^{x_{1}}\delta_{1}^{x_{3}}+\delta_{1}^{x_{1}}\delta_{0}^{x_{3}}+\delta_{1}^{x_{1}}\delta_{1}^{x_{3}}\right],\\ p(x_{2}x_{3}|X_{2}X_{3})&=\frac{1}{4}\left[\delta_{0}^{x_{2}}\delta_{0}^{x_{3}}+\delta_{0}^{x_{2}}\delta_{1}^{x_{3}}+\delta_{1}^{x_{2}}\delta_{0}^{x_{3}}+\delta_{1}^{x_{2}}\delta_{1}^{x_{3}}\right],\end{split}$ (21)
where $\delta_{a}^{b}$ is the Kronecker delta. Since all bipartite marginals are maximally mixed and the contextual fraction is $\mathrm{CF}=1$, the GHZ correlation constitutes an AMCC.
The three-way nonlocal correlations are given by:
$p(x_{1}x_{2}x_{3}|X_{1}X_{2}X_{3})=\begin{cases}1/4,&\text{if }x_{1}\oplus x_{2}\oplus x_{3}=X_{1}X_{2}X_{3},\\ 0,&\text{otherwise.}\end{cases}$ (22)
These correlations are maximally mixed, as can be verified by direct calculation of the marginals. Using the linear optimization method described earlier, we computed the contextual fraction (CF) for this distribution and obtained $\mathrm{CF}=1$, confirming that it is an Absolutely Maximally Contextual Correlation (AMCC). We also verified that the other three-way nonlocal correlations listed in Ref. Barrett et al. (2005) are likewise AMCCs.
There exist cases in the $(3,2,2)$-Bell scenario that are maximally contextual but do not qualify as AMCCs, as they fail to satisfy the maximal marginals condition. For instance, consider the correlations presented in Table 1: their bipartite marginals are not uniformly random.
Con.& Sec. $(0,0,0)$ $(0,0,1)$ $(0,1,0)$ $(0,1,1)$ $(1,0,0)$ $(1,0,1)$ $(1,1,0)$ $(1,1,1)$
$(0,0,0)$ 0 $\tfrac{1}{4}$ $\tfrac{1}{2}$ 0 $\tfrac{1}{4}$ 0 0 0
$(0,0,1)$ $\tfrac{1}{4}$ 0 0 $\tfrac{1}{2}$ 0 $\tfrac{1}{4}$ 0 0
$(0,1,0)$ $\tfrac{1}{4}$ 0 $\tfrac{1}{4}$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ 0 0 0
$(0,1,1)$ $\tfrac{1}{4}$ 0 0 $\tfrac{1}{2}$ 0 $\tfrac{1}{4}$ 0 0
$(1,0,0)$ 0 $\tfrac{1}{4}$ $\tfrac{1}{2}$ 0 $\tfrac{1}{4}$ 0 0 0
$(1,0,1)$ $\tfrac{1}{4}$ 0 0 $\tfrac{1}{2}$ 0 $\tfrac{1}{4}$ 0 0
$(1,1,0)$ $\tfrac{1}{2}$ 0 0 $\tfrac{1}{4}$ 0 0 $\tfrac{1}{4}$ 0
$(1,1,1)$ $\tfrac{1}{4}$ $\tfrac{1}{4}$ 0 $\tfrac{1}{4}$ 0 0 0 $\tfrac{1}{4}$
Table 1: Strongly contextual table for (3,2,2)- scenario but not AMCC.
We present additional instances of AMCC for the $(3,2,2)$ scenario here. We provide a general construction and will extend it to broader families in the next section. We begin by considering a parametrized family of no-signaling $(3,2,2)$ Bell scenarios that satisfy the normalization condition. The corresponding polytope has dimension 26 (reflecting 26 independent parameters), as detailed in Appendix (B). After further parameter reduction, the resulting parametric table for AMCCs is shown in Table 2.
Con.& Sec. $(0,0,0)$ $(0,0,1)$ $(0,1,0)$ $(0,1,1)$ $(1,0,0)$ $(1,0,1)$ $(1,1,0)$ $(1,1,1)$
$(0,0,0)$ $p_{1}$ $\tfrac{1}{4}-p_{1}$ $\tfrac{1}{4}-p_{1}$ $p_{1}$ $\tfrac{1}{4}-p_{1}$ $p_{1}$ $p_{1}$ $\tfrac{1}{4}-p_{1}$
$(0,0,1)$ $p_{2}$ $\tfrac{1}{4}-p_{2}$ $\tfrac{1}{4}-p_{2}$ $p_{2}$ $\tfrac{1}{4}-p_{2}$ $p_{2}$ $p_{2}$ $\tfrac{1}{4}-p_{2}$
$(0,1,0)$ $p_{3}$ $\tfrac{1}{4}-p_{3}$ $\tfrac{1}{4}-p_{3}$ $p_{3}$ $\tfrac{1}{4}-p_{3}$ $p_{3}$ $p_{3}$ $\tfrac{1}{4}-p_{3}$
$(0,1,1)$ $p_{4}$ $\tfrac{1}{4}-p_{4}$ $\tfrac{1}{4}-p_{4}$ $p_{4}$ $\tfrac{1}{4}-p_{4}$ $p_{4}$ $p_{4}$ $\tfrac{1}{4}-p_{4}$
$(1,0,0)$ $p_{5}$ $\tfrac{1}{4}-p_{5}$ $\tfrac{1}{4}-p_{5}$ $p_{5}$ $\tfrac{1}{4}-p_{5}$ $p_{5}$ $p_{5}$ $\tfrac{1}{4}-p_{5}$
$(1,0,1)$ $p_{6}$ $\tfrac{1}{4}-p_{6}$ $\tfrac{1}{4}-p_{6}$ $p_{6}$ $\tfrac{1}{4}-p_{6}$ $p_{6}$ $p_{6}$ $\tfrac{1}{4}-p_{6}$
$(1,1,0)$ $p_{7}$ $\tfrac{1}{4}-p_{7}$ $\tfrac{1}{4}-p_{7}$ $p_{7}$ $\tfrac{1}{4}-p_{7}$ $p_{7}$ $p_{7}$ $\tfrac{1}{4}-p_{7}$
$(1,1,1)$ $p_{8}$ $\tfrac{1}{4}-p_{8}$ $\tfrac{1}{4}-p_{8}$ $p_{8}$ $\tfrac{1}{4}-p_{8}$ $p_{8}$ $p_{8}$ $\tfrac{1}{4}-p_{8}$
Table 2: Reduced (8 free parameters) probability table.
Table 2 now contains only $8$ free parameters. These parameters can only take values in $[0,1/4]$ to preserve the non-negativity of the probabilities. The empirical model corresponding to this table has maximal marginals for all parameter choices.
We examine which conditions on these parameters yield a contextual fraction (CF) of $1$, using the optimization method described earlier. When we fix two of the eight parameters to $0.125$ or $0.25$, and set the remaining parameters to 0, the possible assignments lead to exactly three possible CF values: $\mathrm{CF}\in\{0,0.5,1\}$. These are not the only parameter choices that yield a maximally contextual model. Our results show that if one parameter is set to $0.25$, the remaining seven can take any values in $[0,0.125]$, and the CF remains $1$. Moreover, if two parameters are fixed to any values (including zero) within the allowed range, the CF is always $1$. This demonstrates that the eight-parameter empirical model shown in Table 2, with the parameter combinations described above, yields $\mathrm{CF}=1$ and the maximal marginals by construction. Hence, all such models correspond to AMCCs. We generalize these results in the next section using the Constraint Satisfaction Problem (CSP) approach, employing Table 2 to construct the Boolean empirical model.
IV The Construction of AMC and non-AMC Correlations
Since the evidence provided in the previous section indicates that there exist maximally contextual correlations (MCCs) that are not AMCCs (see Table 1), the relation is asymmetric. Moreover, from the parametric Table 2, we have observed several examples of AMCCs by conditioning on the parameters. In this section, we present systematic constructions that generate both AMCC and non-AMCC  [^2] The correlation with the contextual fraction $\mathrm{CF}=1$ but not the maximal marginal models in a unified framework. We also provide proofs for certain properties of the empirical models introduced in the previous section.
IV.1 CSP Model
To search for new classes of AMCCs, we must first understand the structural properties of correlations. A useful approach is to examine the possibilistic form of correlations, also known as logical contextual models Santos and Amaral (2021). Given a scenario $\langle\mathcal{X},\mathcal{M},\mathcal{O}\rangle$, a distribution over the Boolean semiring $\mathbb{B}=(\{0,1\},\vee,\wedge,\neg)$ is called a possibilistic distribution. That is, a mapping $\bar{\mathcal{P}}:\mathcal{E}\to\mathbb{B}$ with $\sum_{x\in\mathcal{E}}\bar{\mathcal{P}}(x)=1$ defines an element of $D_{\mathbb{B}}(\mathcal{E})$. For any probabilistic empirical model, the corresponding possibilistic collapse is obtained by taking the support:
$\bar{e}_{C}(s)=\begin{cases}1,&\text{if }e_{C}(s)>0,\\ 0,&\text{otherwise},\end{cases}\qquad\forall C\in\mathcal{M}.$ (23)
The set of global sections compatible with the support of a probabilistic empirical model $e=\{e_{C}\}_{C\in\mathcal{M}}$ is
$S_{e}=\{\,s\in\mathcal{E}(\mathcal{X})\mid\forall C\in\mathcal{M},\;s|_{C}\in\operatorname{supp}(e_{C})\,\}.$ (24)
The model $e$ is strongly contextual if and only if $S_{e}=\varnothing$. Equivalently, there exists no global distribution $d\in D_{\mathbb{B}}(\mathcal{X})$ such that $d|_{C}\in\operatorname{supp}(e_{C})$ for all $C\in\mathcal{M}$. From Definition (6), strong contextuality coincides with maximal contextuality.
This characterization can be mapped to a Constraint Satisfaction Problem (CSP), which for dichotomic measurements reduces to a Boolean satisfiability problem (SAT). Given an empirical model $e$, we associate a set of Boolean variables $\bar{\mathcal{X}}$ corresponding to the observables in the probabilistic model. For each context $C\in\mathcal{M}$ and each section $s\in S_{e}(C)$ in the support, we define a Boolean proposition. The overall propositional formula for the model is
$B_{e}\;:=\;\bigwedge_{C\in\mathcal{M}}\Bigl(\;\bigvee_{s\in S_{e}(C)}b_{s}\;\Bigr),$ (25)
where each $b_{s}$ is a Boolean statement encoding the assignment prescribed by the section $s$:
$b_{s}\;:=\;\Bigl(\;\bigwedge_{\begin{subarray}{c}X\in C\\ s(X)=1\end{subarray}}X\;\Bigr)\;\wedge\;\Bigl(\;\bigwedge_{\begin{subarray}{c}X\in C\\ s(X)=0\end{subarray}}\neg X\;\Bigr).$ (26)
Thus $B_{e}$ is a conjunction over contexts of the disjunctions of all supported local assignments.
As an illustration, consider the $(2,2,2)$ PR‑box empirical model Eq.(18). Its possibilistic collapse is
$\begin{array}[]{c|cccc}X_{1}X_{2}\backslash x_{1}x_{2}&(0,0)&(0,1)&(1,0)&(1,1)\\ \hline\cr(0,0)&1&0&0&1\\ (0,1)&1&0&0&1\\ (1,0)&1&0&0&1\\ (1,1)&0&1&1&0\end{array}$ (27)
which yields the Boolean formulas
$B_{1}$ $=(X_{1}\wedge X_{2})\;\vee\;(\neg X_{1}\wedge\neg X_{2}),$ (28)
$B_{2}$ $=(X_{1}\wedge X_{2}^{\prime})\;\vee\;(\neg X_{1}\wedge\neg X_{2}^{\prime}),$
$B_{3}$ $=(X_{1}^{\prime}\wedge X_{2})\;\vee\;(\neg X_{1}^{\prime}\wedge\neg X_{2}),$
$B_{4}$ $=(\neg X_{1}^{\prime}\wedge X_{2}^{\prime})\;\vee\;(X_{1}^{\prime}\wedge\neg X_{2}^{\prime}).$
The first three formulas impose correlations ($x_{1}=x_{2}$, $x_{1}=x_{2}^{\prime}$, $x_{1}^{\prime}=x_{2}$), while the fourth imposes an anti‑correlation ($x_{1}^{\prime}\neq x_{2}^{\prime}$). This set of conditions is logically inconsistent: the first three together imply $x_{1}^{\prime}=x_{2}^{\prime}$, which contradicts the fourth condition. Thus, no assignment in the set $\{(x_{1},x_{2},x_{1}^{\prime},x_{2}^{\prime})\mid x_{1},x_{2},x_{1}^{\prime},x_{2}^{\prime}\in\{0,1\}\}$ simultaneously satisfies all four Boolean formulas, implying $S_{e}=\varnothing$. Hence, the model is strongly contextual and has $\mathrm{CF}=1$.
We now introduce a new family of constructions that produce only AMCCs and are symmetric in structure. By symmetric, we mean that in each row of the probability table, exactly four entries are 0 and the remaining four are $1/4$.
IV.2 Parity Check method
This mathematical construction is based on Mermin’s proof and generates only symmetric AMCCs. In this method, we consider a set of equations, each involving a modulo 2 sum of observables. Each equation takes the form $\sum X=P\;(\mathrm{mod}\;2)$, where $X\in\mathcal{X}$ and the parity $P\in\{0,1\}$. The parity is even for $P=0$ and odd for $P=1$. For example, in the $(2,2,2)$ scenario, the equations are:
$X_{1}+X_{2}$ $=P_{1}\quad(\mathrm{mod}2),$ (29)
$X_{1}+X_{2}^{\prime}$ $=P_{2}\quad(\mathrm{mod}2),$
$X_{1}^{\prime}+X_{2}$ $=P_{3}\quad(\mathrm{mod}2),$
$X_{1}^{\prime}+X_{2}^{\prime}$ $=P_{4}\quad(\mathrm{mod}2).$
There are $2^{4}=16$ possible combinations of parity choices $(P_{1},P_{2},P_{3},P_{4})\in\{0,1\}^{4}$. The Boolean statements $b_{s}$ are selected according to the parity of each equation. Each assignment yields a corresponding set of Boolean formulas $B_{e}$.
Consider the specific choices $P_{1}=P_{2}=P_{3}=0$ and $P_{4}=1$:
$X_{1}+X_{2}$ $=0\quad(\mathrm{mod}2),$ (30)
$X_{1}+X_{2}^{\prime}$ $=0\quad(\mathrm{mod}2),$
$X_{1}^{\prime}+X_{2}$ $=0\quad(\mathrm{mod}2),$
$X_{1}^{\prime}+X_{2}^{\prime}$ $=1\quad(\mathrm{mod}2).$
Upon summing up all the equations, we obtain
$2X_{1}+2X_{2}+2X_{1}^{\prime}+2X_{2}^{\prime}\equiv 1\pmod{2}$, (31)
which simplifies to $0\equiv 1$, a contradiction. Therefore, it follows that no assignment can simultaneously satisfy all four equations. To determine the corresponding Boolean formulas, the choice of boolean statements are distinguished into two sets: for even parity ($P=0$) the allowed choices are $\{(\neg X_{1}\wedge\neg X_{2}),(X_{1}\wedge X_{2})\}$, while for odd parity ($P=1$) they are $\{(\neg X_{1}\wedge X_{2}),(X_{1}\wedge\neg X_{2})\}$. This set of choices is the same for $X_{1}^{\prime}$ and $X_{2}^{\prime}$. Here, the boolean formula corresponding to $X_{1}+X_{2}=0\pmod{2}$ is
$B_{1}=(X_{1}\wedge X_{2})\;\vee\;(\neg X_{1}\wedge\neg X_{2}).$
This mapping is consistent since, for binary valuation, the modulo 2 operation and the logical OR operation are identical. Here, this parity equation is satisfiable only when both variables are the same. After mapping to the Boolean form, $B_{1}$ is satisfiable when both Boolean variables are the same. Similarly, the other Boolean formulas, $B_{1}$, $B_{2}$, and $B_{3}$, can be obtained. The resulting set of Boolean formulas corresponds to Eq. (28), which describes a strongly contextual model and an AMCC for the $(2,2,2)$ scenario. The logical contradiction persists from the parity equation to the Boolean formula mapping because the reduced equation Eq. (31) and the Boolean equation Eq. (25) are identical. For instance,
$B_{e}=B_{1}\;\wedge\;B_{2}\;\wedge\;B_{3}\;\wedge\;B_{4}$ (32)
If any of the Boolean formulas are false, then $B_{e}$ will be False. Namely, the first parity equation is satisfied by only $2$ assignments $(0,0)$ and $(1,1)$ out of $4$ possible assignments, and the corresponding Boolean formula $B_{1}$ does not satisfy the last parity equation, as well as the corresponding Boolean formula $B_{4}$. Hence, the reduced parity equation Eq.(31) shows the contradiction as well as the corresponding Boolean equation Eq.(32). Accordingly, if no assignment $(x1,x2,x1^{\prime},x2^{\prime})\in\{0,1\}^{4}$ satisfies all parity equations simultaneously, the associated Boolean formulas are unsatisfiable. Consequently, the empirical model exhibits strong contextuality. The possibilistic model derived from the parity equations corresponds to a possibilistic collapse of a probabilistic model, which can be constructed from the parametric table by setting the parameters corresponding to absent terms in the Boolean statements to zero, i.e., the zero constraints.
The reduced parametric table for the $(2,2,2)$ scenario contains $8$ free parameters, as determined by the normalization and no-signaling constraints. Maximal marginality is ensured by the symmetry of the possibilistic model: each Boolean formula contains exactly half the total number of Boolean statements. Therefore, to satisfy the normalization condition and no-signaling condition, all parameters within each row for bipartite marginals must be equal, which ensures that the correlations are maximally marginal.
Since the model is strongly contextual, the corresponding probabilistic model is also maximally contextual. Furthermore, maximal marginality is ensured by the symmetric structure of the possibilistic model. Thus, the resulting correlations constitute an Absolutely Maximally Contextual Correlation (AMCC).
For the $(3,2,2)$ scenario, let $X\in\{X_{1},X_{2},X_{3},X_{1}^{\prime},X_{2}^{\prime},X_{3}^{\prime}\}$ with each $X\in\{0,1\}$. The parity equations are:
$X_{1}+X_{2}+X_{3}$ $=P_{1}\quad(\mathrm{mod}2),$ (33)
$X_{1}+X_{2}+X_{3}^{\prime}$ $=P_{2}\quad(\mathrm{mod}2),$
$X_{1}+X_{2}^{\prime}+X_{3}$ $=P_{3}\quad(\mathrm{mod}2),$
$X_{1}+X_{2}^{\prime}+X_{3}^{\prime}$ $=P_{4}\quad(\mathrm{mod}2),$
$X_{1}^{\prime}+X_{2}+X_{3}$ $=P_{5}\quad(\mathrm{mod}2),$
$X_{1}^{\prime}+X_{2}+X_{3}^{\prime}$ $=P_{6}\quad(\mathrm{mod}2),$
$X_{1}^{\prime}+X_{2}^{\prime}+X_{3}$ $=P_{7}\quad(\mathrm{mod}2),$
$X_{1}^{\prime}+X_{2}^{\prime}+X_{3}^{\prime}$ $=P_{8}\quad(\mathrm{mod}2).$
[…]
$X_{1}+X_{2}+X_{3}$ $=0\quad(\mathrm{mod}2),$ (34)
$X_{1}+X_{2}+X_{3}^{\prime}$ $=1\quad(\mathrm{mod}2),$
$X_{1}+X_{2}^{\prime}+X_{3}$ $=1\quad(\mathrm{mod}2),$
$X_{1}+X_{2}^{\prime}+X_{3}^{\prime}$ $=1\quad(\mathrm{mod}2),$
$X_{1}^{\prime}+X_{2}+X_{3}$ $=1\quad(\mathrm{mod}2),$
$X_{1}^{\prime}+X_{2}+X_{3}^{\prime}$ $=1\quad(\mathrm{mod}2),$
$X_{1}^{\prime}+X_{2}^{\prime}+X_{3}$ $=1\quad(\mathrm{mod}2),$
$X_{1}^{\prime}+X_{2}^{\prime}+X_{3}^{\prime}$ $=1\quad(\mathrm{mod}2).$
Summing all eight equations modulo 2 gives
$4X_{1}+4X_{1}^{\prime}+4X_{2}+4X_{2}^{\prime}+4X_{3}+4X_{3}^{\prime}\equiv 7\pmod{2}$, (35)
which reduces to $0\equiv 1\pmod{2}$, a logical contradiction. The corresponding Boolean equation
$B_{e}=B_{1}\;\wedge\cdots\wedge\;B_{8},$ (36)
where $B_{1},\cdots,B_{8}$ are all boolean formulas constructed from Eq.(34). Namely, the first parity equation
$X_{1}+X_{2}+X_{3}=0\pmod{2}$,
can be mapped to the Boolean formula
$B_{1}=$ $(\neg X_{1}\wedge\neg X_{2}\wedge\neg X_{3})\;\vee\;(\neg X_{1}\wedge X_{2}\wedge X_{3})$
$\vee\;(X_{1}\wedge\neg X_{2}\wedge X_{3})\;\vee\;(X_{1}\wedge X_{2}\wedge\neg X_{3})\,.$
The construction relies on the fact that the parity equation is satisfied by only $4$ assignments $(0,0,0),(0,1,1),(1,0,1)$ and $(1,1,0)$ out of a total of $8$ possible assignments. Accordingly, the corresponding Boolean formula is satisfiable only for these assignments. The persistence of the contradiction from the parity equations to the Boolean formulas can be demonstrated as follows. Consider the first and second parity equations. The second parity equation,
$X1+X2+X3^{\prime}=1\pmod{2}$,
implies that $(X1+X2=X3^{\prime}\pmod{2})$. Combining this with the first parity equation yields $(X3=X3^{\prime}\pmod{2})$, which is a contradiction. Specifically, for the assignments that satisfy the first parity equation and, thus, the corresponding Boolean formula, the second parity equation, together with its associated Boolean formula, becomes unsatisfiable. Hence, this leads to the contradiction shown in Eq. (35), which, in turn, results in the unsatisfiability of the corresponding Boolean equation, Eq. (36). Hence, no consistent assignment exists, i.e., $S_{e}=\varnothing$ for the possibilistic model. Therefore, the empirical model is strongly contextual, and maximal marginality is ensured by the symmetric structure of zero constraints. Specifically, since half of the total Boolean statements are selected for each row, assigning zeros to the absent Boolean statements in the parametric table 4 results in only four nonzero elements per row. To satisfy the normalization condition and the no-signaling constraint, it follows that all nonzero parameters within a row must be equal, which ensures that the bipartite marginals are maximal. Consequently, the correlations associated with the parity equations constitute an AMCC.
IV.3 Construction of AMCC and non-AMCC
In this section, we construct both AMCC and non‑AMCC models using the parity‑check method and the CSP approach. We begin with the parity‑check construction. Depending on the chosen parity values, we select the corresponding Boolean statements and assemble the resulting Boolean formula. We then test whether the set of parity equations is simultaneously unsatisfiable for all possible assignments. If the associated Boolean formulas are unsatisfiable for every assignment, the corresponding possibilistic model is strongly contextual.
Numerical analysis shows that, out of the $256$ possible parity assignments, exactly $240$ give rise to AMCCs. As an example, consider the parity choice $P_{1}=P_{2}=P_{3}=P_{4}=P_{5}=P_{6}=P_{7}=0,\qquad P_{8}=1.$ The corresponding Boolean formula is
$B_{1}$ $=(\neg X_{1}\wedge\neg X_{2}\wedge\neg X_{3})\;\vee\;(\neg X_{1}\wedge X_{2}\wedge X_{3})\;\vee\;(X_{1}\wedge\neg X_{2}\wedge X_{3})\;\vee\;(X_{1}\wedge X_{2}\wedge\neg X_{3})\;,$ (37)
$B_{2}$ $=(\neg X_{1}\wedge\neg X_{2}\wedge\neg X_{3}^{\prime})\;\vee\;(\neg X_{1}\wedge X_{2}\wedge X_{3}^{\prime})\;\vee\;(X_{1}\wedge\neg X_{2}\wedge X_{3}^{\prime})\;\vee\;(X_{1}\wedge X_{2}\wedge\neg X_{3}^{\prime})\;,$
$B_{3}$ $=(\neg X_{1}\wedge\neg X_{2}^{\prime}\wedge\neg X_{3})\;\vee\;(\neg X_{1}\wedge X_{2}^{\prime}\wedge X_{3})\;\vee\;(X_{1}\wedge\neg X_{2}^{\prime}\wedge X_{3})\;\vee\;(X_{1}\wedge X_{2}^{\prime}\wedge\neg X_{3})\;,$
$B_{4}$ $=(\neg X_{1}\wedge\neg X_{2}^{\prime}\wedge\neg X_{3}^{\prime})\;\vee\;(\neg X_{1}\wedge X_{2}^{\prime}\wedge X_{3}^{\prime})\;\vee\;(X_{1}\wedge\neg X_{2}^{\prime}\wedge X_{3}^{\prime})\;\vee\;(X_{1}\wedge X_{2}^{\prime}\wedge\neg X_{3}^{\prime})\;,$
$B_{5}$ $=(\neg X_{1}^{\prime}\wedge\neg X_{2}\wedge\neg X_{3})\;\vee\;(\neg X_{1}^{\prime}\wedge X_{2}\wedge X_{3})\;\vee\;(X_{1}^{\prime}\wedge\neg X_{2}\wedge X_{3})\;\vee\;(X_{1}^{\prime}\wedge X_{2}\wedge\neg X_{3})\;,$
$B_{6}$ $=(\neg X_{1}^{\prime}\wedge\neg X_{2}\wedge\neg X_{3}^{\prime})\;\vee\;(\neg X_{1}^{\prime}\wedge X_{2}\wedge X_{3}^{\prime})\;\vee\;(X_{1}^{\prime}\wedge\neg X_{2}\wedge X_{3}^{\prime})\;\vee\;(X_{1}^{\prime}\wedge X_{2}\wedge\neg X_{3}^{\prime})\;,$
$B_{7}$ $=(\neg X_{1}^{\prime}\wedge\neg X_{2}^{\prime}\wedge\neg X_{3})\;\vee\;(\neg X_{1}^{\prime}\wedge X_{2}^{\prime}\wedge X_{3})\;\vee\;(X_{1}^{\prime}\wedge\neg X_{2}^{\prime}\wedge X_{3})\;\vee\;(X_{1}^{\prime}\wedge X_{2}^{\prime}\wedge\neg X_{3})\;,$
$B_{8}$ $=(\neg X_{1}^{\prime}\wedge\neg X_{2}^{\prime}\wedge X_{3}^{\prime})\;\vee\;(\neg X_{1}^{\prime}\wedge X_{2}^{\prime}\wedge\neg X_{3}^{\prime})\;\vee\;(X_{1}^{\prime}\wedge\neg X_{2}^{\prime}\wedge\neg X_{3}^{\prime})\;\vee\;(X_{1}^{\prime}\wedge X_{2}^{\prime}\wedge X_{3}^{\prime})\;,$
This is the possibilistic collapse of a probability table whose nonzero entries are all $1/4$, each corresponding to a Boolean statement $b_{s}$ appearing in the Boolean formula. Hence, the parity‑check method provides a systematic and convenient way to generate AMCCs for $(3,2,2)$ models. The models produced by this method have symmetric probability tables, where exactly half of the assignments in each context receive probability $1/4$ and the other half receive probability 0.
In contrast, Tables (1) is asymmetric. We now turn to non‑AMCC and asymmetric empirical models, which, under appropriate conditions, lead to the asymmetric AMCC. To construct such models, we adopt a more general approach by freely selecting Boolean statements $b_{s}$. If we consider all possible combinations of Boolean statements across the eight contexts, there are $255^{8}$ possibilities (since each context can contain any non‑empty subset of its $8$ possible local assignments). Previously, the parity condition provided a simple rule for selecting statements, and the number of allowed combinations was small. Now, scanning all combinations by brute force is intractable.
To overcome this difficulty, we start by fixing one Boolean statement in each context and then systematically vary the remaining choices. In particular, we work with the following set of Boolean formulas (constraints):
$B_{1}$ $=(\neg X_{1}\wedge\neg X_{2}\wedge X_{3})\;\vee\;(\neg X_{1}\wedge X_{2}\wedge\neg X_{3})\;\vee\;(X_{1}\wedge\neg X_{2}\wedge\neg X_{3})\;\vee\;(X_{1}\wedge X_{2}\wedge X_{3})\;,$ (38)
$B_{2}$ $=(\neg X_{1}\wedge\neg X_{2}\wedge\neg X_{3}^{\prime})\;\vee\;(\neg X_{1}\wedge\neg X_{2}\wedge X_{3}^{\prime})\;\vee\;(\neg X_{1}\wedge X_{2}\wedge\neg X_{3}^{\prime})\;\vee\;(\neg X_{1}\wedge X_{2}\wedge X_{3}^{\prime})\;,$
$B_{3}$ $=(\neg X_{1}\wedge\neg X_{2}^{\prime}\wedge\neg X_{3})\;\vee\;(\neg X_{1}\wedge\neg X_{2}^{\prime}\wedge X_{3})\;\vee\;(\neg X_{1}\wedge X_{2}^{\prime}\wedge\neg X_{3})\;\vee\;(\neg X_{1}\wedge X_{2}^{\prime}\wedge X_{3})\;,$
$B_{4}$ $=(\neg X_{1}\wedge\neg X_{2}^{\prime}\wedge\neg X_{3}^{\prime})\;\vee\;(\neg X_{1}\wedge X_{2}^{\prime}\wedge X_{3}^{\prime})\;\vee\;(X_{1}\wedge\neg X_{2}^{\prime}\wedge X_{3}^{\prime})\;\vee\;(X_{1}\wedge X_{2}^{\prime}\wedge\neg X_{3}^{\prime})\;,$
$B_{5}$ $=(\neg X_{1}^{\prime}\wedge\neg X_{2}\wedge\neg X_{3})\;\vee\;(\neg X_{1}^{\prime}\wedge\neg X_{2}\wedge X_{3})\;\vee\;(\neg X_{1}^{\prime}\wedge X_{2}\wedge\neg X_{3})\;\vee\;(\neg X_{1}^{\prime}\wedge X_{2}\wedge X_{3})\;,$
$B_{6}$ $=(\neg X_{1}^{\prime}\wedge\neg X_{2}\wedge\neg X_{3}^{\prime})\;\vee\;(\neg X_{1}^{\prime}\wedge X_{2}\wedge X_{3}^{\prime})\;\vee\;(X_{1}^{\prime}\wedge\neg X_{2}\wedge X_{3}^{\prime})\;\vee\;(X_{1}^{\prime}\wedge X_{2}\wedge\neg X_{3}^{\prime})\;,$
$B_{7}$ $=(\neg X_{1}^{\prime}\wedge\neg X_{2}^{\prime}\wedge\neg X_{3})\;\vee\;(\neg X_{1}^{\prime}\wedge X_{2}^{\prime}\wedge X_{3})\;\vee\;(X_{1}^{\prime}\wedge\neg X_{2}^{\prime}\wedge X_{3})\;\vee\;(X_{1}^{\prime}\wedge X_{2}^{\prime}\wedge\neg X_{3})\;,$
$B_{8}$ $=(\neg X_{1}^{\prime}\wedge\neg X_{2}^{\prime}\wedge\neg X_{3}^{\prime})\;\vee\;(\neg X_{1}^{\prime}\wedge\neg X_{2}^{\prime}\wedge X_{3}^{\prime})\;\vee\;(\neg X_{1}^{\prime}\wedge X_{2}^{\prime}\wedge\neg X_{3}^{\prime})\;\vee\;(\neg X_{1}^{\prime}\wedge X_{2}^{\prime}\wedge X_{3}^{\prime})\;,$
One could in principle start with any set of admissible Boolean formulas; we choose the set above for convenience. For each Boolean formula, there are four remaining terms (local assignments) that can be added. Considering all combinations of these extra terms would yield approximately $4.2\times 10^{9}$ possibilities, which is still too large to handle exhaustively. We therefore reduce the complexity by only adding terms that are absent from the formulas $B_{2},B_{3},B_{5},B_{8}$. With this restriction, the number of combinations drops to $65,536$. Among these, only $2,401$ Boolean formulas satisfy the Boolean no‑signaling condition (Appendix (C)) and are simultaneously unsatisfiable as CSP instances. Consequently, all $2,401$ such sets of Boolean formulas correspond to strongly contextual empirical models. One of them is as follows;
$B_{1}$ $=(\neg X_{1}\wedge\neg X_{2}\wedge X_{3})\vee(\neg X_{1}\wedge X_{2}\wedge\neg X_{3})\vee(X_{1}\wedge\neg X_{2}\wedge\neg X_{3})\vee(X_{1}\wedge X_{2}\wedge X_{3}),$ (39)
$B_{2}$ $=(\neg X_{1}\wedge\neg X_{2}\wedge\neg X_{3}^{\prime})\vee(\neg X_{1}\wedge\neg X_{2}\wedge X_{3}^{\prime})\vee(\neg X_{1}\wedge X_{2}\wedge\neg X_{3}^{\prime})\vee(\neg X_{1}\wedge X_{2}\wedge X_{3}^{\prime})$
$\vee(X_{1}\wedge\neg X_{2}\wedge\neg X_{3}^{\prime})\vee(X_{1}\wedge X_{2}\wedge X_{3}^{\prime}),$
$B_{3}$ $=(\neg X_{1}\wedge\neg X_{2}^{\prime}\wedge\neg X_{3})\vee(\neg X_{1}\wedge\neg X_{2}^{\prime}\wedge X_{3})\vee(\neg X_{1}\wedge X_{2}^{\prime}\wedge\neg X_{3})\vee(\neg X_{1}\wedge X_{2}^{\prime}\wedge X_{3})$
$\vee(X_{1}\wedge\neg X_{2}^{\prime}\wedge\neg X_{3})\vee(X_{1}\wedge X_{2}^{\prime}\wedge X_{3}),$
$B_{4}$ $=(\neg X_{1}\wedge\neg X_{2}^{\prime}\wedge\neg X_{3}^{\prime})\vee(\neg X_{1}\wedge X_{2}^{\prime}\wedge X_{3}^{\prime})\vee(X_{1}\wedge\neg X_{2}^{\prime}\wedge X_{3}^{\prime})\vee(X_{1}\wedge X_{2}^{\prime}\wedge\neg X_{3}^{\prime}),$
$B_{5}$ $=(\neg X_{1}^{\prime}\wedge\neg X_{2}\wedge\neg X_{3})\vee(\neg X_{1}^{\prime}\wedge\neg X_{2}\wedge X_{3})\vee(\neg X_{1}^{\prime}\wedge X_{2}\wedge\neg X_{3})\vee(\neg X_{1}^{\prime}\wedge X_{2}\wedge X_{3})$
$\vee(X_{1}^{\prime}\wedge\neg X_{2}\wedge\neg X_{3})\vee(X_{1}^{\prime}\wedge X_{2}\wedge X_{3}),$
$B_{6}$ $=(\neg X_{1}^{\prime}\wedge\neg X_{2}\wedge\neg X_{3}^{\prime})\vee(\neg X_{1}^{\prime}\wedge X_{2}\wedge X_{3}^{\prime})\vee(X_{1}^{\prime}\wedge\neg X_{2}\wedge X_{3}^{\prime})\vee(X_{1}^{\prime}\wedge X_{2}\wedge\neg X_{3}^{\prime}),$
$B_{7}$ $=(\neg X_{1}^{\prime}\wedge\neg X_{2}^{\prime}\wedge\neg X_{3})\vee(\neg X_{1}^{\prime}\wedge X_{2}^{\prime}\wedge X_{3})\vee(X_{1}^{\prime}\wedge\neg X_{2}^{\prime}\wedge X_{3})\vee(X_{1}^{\prime}\wedge X_{2}^{\prime}\wedge\neg X_{3}),$
$B_{8}$ $=(\neg X_{1}^{\prime}\wedge\neg X_{2}^{\prime}\wedge\neg X_{3}^{\prime})\vee(\neg X_{1}^{\prime}\wedge\neg X_{2}^{\prime}\wedge X_{3}^{\prime})\vee(\neg X_{1}^{\prime}\wedge X_{2}^{\prime}\wedge\neg X_{3}^{\prime})\vee(\neg X_{1}^{\prime}\wedge X_{2}^{\prime}\wedge X_{3}^{\prime})$
$\vee(X_{1}^{\prime}\wedge\neg X_{2}^{\prime}\wedge\neg X_{3}^{\prime})\vee(X_{1}^{\prime}\wedge X_{2}^{\prime}\wedge X_{3}^{\prime}),$
This set of Boolean formulas corresponds to the collapse of the probabilistic table(3). This reduces the parameters of the table(4) to $3$.
Con.& Sec. (0,0,0) (0,0,1) (0,1,0) (0,1,1) (1,0,0) (1,0,1) (1,1,0) (1,1,1)
$(0,0,0)$ 0 $p_{1}$ $p_{1}$ 0 $1/2-p_{1}$ 0 0 $1/2-p_{1}$
$(0,0,1)$ $p_{2}$ $p_{1}-p_{2}$ $p_{3}$ $p_{1}-p_{3}$ $1/2-p_{1}$ 0 0 $1/2-p_{1}$
$(0,1,0)$ $p_{2}$ $p_{3}$ $p_{1}-p_{2}$ $p_{1}-p_{3}$ $1/2-p_{1}$ 0 0 $1/2-p_{1}$
$(0,1,1)$ $p_{2}+p_{3}$ 0 0 $2p_{1}-p_{2}-p_{3}$ 0 $1/2-p_{1}$ $1/2-p_{1}$ 0
$(1,0,0)$ $1/2-2p_{1}+p_{2}$ $p_{1}$ $p_{1}$ $1/2-p_{1}-p_{3}$ $p_{2}-p_{2}$ 0 0 $p_{3}$
$(1,0,1)$ $1/2-p_{1}+p_{2}$ 0 0 $1/2-p_{3}$ 0 $p_{1}-p_{2}$ $p_{3}$ 0
$(1,1,0)$ $1/2-p_{1}+p_{2}$ 0 0 $1/2-p_{3}$ 0 $p_{3}$ $p_{1}-p_{2}$ 0
$(1,1,1)$ $p_{2}$ $1/2-p_{1}$ $1/2-p_{1}$ $p_{1}-p_{3}$ $p_{3}$ 0 0 $p_{1}-p_{3}$
Table 3: Maximum contextual table $(3,2,2)$-scenario.
The empirical model corresponding to the probability Table (3) is strongly contextual, provided the parameters satisfy the bounds
$0\leq p_{9}<\tfrac{1}{2},\qquad p_{9}<p_{2}<\tfrac{p_{9}}{2}+\tfrac{1}{4},$
$\qquad 0<p_{11}<\min\!\bigl(p_{2},\ \tfrac{1}{2}-p_{2},2p_{2}-p_{9}\bigr).$
This model is a non‑AMCC because it does not satisfy the maximal‑marginals condition of Definition 7.
The same model becomes an AMCC when $p_{2}=p_{11}=1/4$ and $p_{9}=0$. In the parity‑check method, determining strong contextuality reduces to checking the parity equations, which is computationally straightforward. In contrast, the general CSP‑based method requires verifying both the no‑signaling condition and the satisfiability of the Boolean constraints, adding extra complexity. However, the latter approach yields strongly contextual models with asymmetric probability tables, which are not obtainable from the symmetric parity‑based construction.
V Application
In this section, we will discuss how AMCC is involved in various applications, such as secret sharing schemes and randomness extraction.
V.1 Secrete Sharing
In this section, we describe a hybrid secret‑sharing protocol that combines classical and quantum components Gogioso and Zeng (2019). The HBB scheme Hillery et al. (1999) involves a dealer distributing an entangled quantum state among several parties and then communicating the plaintext via a classical channel, with the actual secret encoded within it. The original secret can only be recovered when all players cooperate.
This scheme can be extended using a generalized Mermin‑type argument, which corresponds precisely to the set of parity equations we employed to generate AMCCs. Suppose a dealer (Alice) wishes to share a secret $S$. She first prepares a plaintext by combining the secret $S$ with a random pad $q$, i.e., $p_{t}=S\oplus q$ — this constitutes the classical part of the protocol. In the quantum part, the resource state is shared so that each player holds $k$ shares, and Alice retains nkn-k shares. Thus, each party can be regarded as possessing $k$ measurement devices; in each round $r$, the inputs are $i_{r}\in\{0,\dots,m\}$ and the outputs are $o_{r}\in P(\bullet)$. The inputs are chosen using independent classical randomness sources, with Alice’s source determining whether a given round is a “secret” round or a “test” round with non‑zero probability. All parties also have access to an authenticated classical channel. To maintain device independence, the quantum resource may be any entangled state.
Alice generates a key only if the protocol succeeds. Her key in round $r$ is obtained from the measurement outcomes on her device as
$a_{\text{dealer}}^{r}=\bigoplus_{j=k+1}^{n}a_{j}^{r}.$ (40)
The players’ inputs are communicated to Alice, who uses them to verify whether the observed correlations form a valid measurement context. This verification checks for perfect correlations that satisfy the parity equations. If the parity conditions are not satisfied, she aborts the protocol. If the correlations are perfect and correspond to a valid context, she treats the round as a test round, shares her ciphertext $q^{r}=p_{t}^{r}\oplus S^{r}$, and also broadcasts the associated phase value $ph^{r}\in\{0,\dots,s\}$ corresponding to the measurement context. By combining their outcomes, the ciphertext, and the phase information, the players can jointly reconstruct Alice’s original secret. The parity conditions employed in the protocol are generalized forms of Mermin-type arguments, meaning that the parity equations cannot be simultaneously satisfied by any global assignment. Consequently, the empirical model associated with the protocol is maximally contextual. As previously discussed, the Boolean formulas corresponding to these parity conditions can be mapped to a probabilistic form through the application of zero constraints. Due to the symmetric structure of the Boolean formulas, the resulting probabilistic model exhibits maximal marginality. Hence, the correlations associated with the parity conditions are AMCCs.
V.2 Randomness extraction
Random number generation is a challenging task both in practice and in mathematical characterization. Random number generators (RNGs) play a crucial role in device‑independent protocols, including cryptographic applications. However, such devices are vulnerable to adversarial manipulation, making it essential to obtain randomness that is certified in a device‑independent (DI) manner. This certification can be achieved by violating a Bell inequality: correlations obtained from the inputs and outputs of a device that violate a Bell inequality cannot be explained by local hidden variables or predetermined strategies. Such a guarantee is therefore called randomness certified by Bell’s theorem Pironio et al. (2010).
The amount of randomness is quantified by the conditional min‑entropy
$H_{\min}(x_{1}x_{2}\mid X_{1}X_{2})=-\log_{2}\max_{x_{1},x_{2}}P(x_{1}x_{2}\mid X_{1}X_{2}).$ (41)
A lower bound for this quantity can be expressed as
$H_{\min}(x_{1}x_{2}\mid X_{1}X_{2})\geq H_{\min}(x_{1}\mid X_{2}),$ (42)
where the local min‑entropy is defined by $H_{\min}(x\mid X)=-\log_{2}\max_{x,X}P(x\mid X)$.
For an AMCC, every reduced bipartition exhibits maximally random correlations. Consequently, an adversary trying to guess the outcomes $x_{1},\dots,x_{k}$ given the inputs $X_{1},\dots,X_{k}$ faces the probability
$p_{\mathrm{guess}}=\max_{x_{1},\dots,x_{k}}P(x_{1},\dots,x_{k}\mid X_{1},\dots,X_{k})=\frac{1}{2^{k}}\qquad\forall\;k<n.$ (43)
The corresponding min‑entropy is therefore
$H_{\min}(x_{1},\dots,x_{k}\mid X_{1},\dots,X_{k})$ $=-\log_{2}p_{\mathrm{guess}}$ (44)
$=\log_{2}2^{k}=k.$
Thus, we obtain the lower bound on the global min‑entropy
$H_{\min}(xy\mid XY)\geq k.$ (45)
This bound shows that the local randomness is maximal: an adversary gains no information about Alice’s outcome and cannot simulate it using any pre‑shared local randomness.
VI Conclusion
Non-local correlations have become an integral part of studies in all aspects of information processing involving quantum resources. While empirical models based on correlations are well-studied for the $(2,2,2)$ scenario, extending these concepts to multipartite systems remains an active area of research. Notably, the characterization of various forms of maximality in the multipartite scenario, similar to that of multipartite maximal entanglement, has remained unclear.
To address this, we introduce a new class of contextual correlations in the multipartite setting, specifically motivated by absolutely maximally entangled (AME) states Helwig et al. (2012) and maximal contextual correlations Liu et al. (2024). These correlations, which we define as AMCCs, lie on the face of a polytope and are both maximally contextual and maximally marginal. The distinction introduced by AMCCs is particularly noteworthy: while AME states for two parties are the Bell states Helwig et al. (2012), the two-party, two-input, two-output AMCCs correspond to PR boxes, which do not have any quantum realization. The three-qubit AME states correspond to three-qubit GHZ states Helwig et al. (2012), and in terms of correlations, the GHZ correlations represent the $(3,2,2)$ AMCCs. Additionally, there exist infinitely many such correlations, as demonstrated in Table (2), which features eight free parameters. Furthermore, although four-qubit AME states do not exist, the parity argument demonstrates that four-party AMCCs are possible. We have also introduced non-AMCC correlations that are maximally contextual but not maximally marginal, which are constructed using the CSP method.
The results presented here can be extended in future work to include the findings of Ref. Aravinda and Srikanth (2013), where the complementarity trade-off between no-signaling and local indeterminacy for quantum correlations in the $(2,2,2)$ scenario is introduced. Their work presents a stronger version of Bell’s theorem and provides a quantitative definition of a bound on randomness. This scope is justified by the fact that the violation of a generalized Bell inequality up to its algebraic bound occurs if and only if the model is strongly contextual Abramsky et al. (2018). The very existence of non-AMCCs, which are maximally contextual but not maximally marginal, provides a departure from the complementarity seen in the bipartite scenario, thereby offering insights into the deeper relationship between nonlocality, randomness, and other nonclassical features in multipartite correlations.
Appendix A Proper faces and AVN
The proper face $F$ Pironio (2005), with dimension $\dim F\leq\dim P-1$, can be defined by imposing zero constraints:
$F=\{\mathbf{p}\in\mathbf{NS}\;|\;p=0\},$ (46)
where $p$ is an element of $\mathbf{p}={p(x_{1},\cdots,x_{i}\mid X_{1},\cdots,X_{i})_{i\leq\dim P-1}})$, and $\mathbf{NS}\subset\mathbb{R}^{r}$ denotes the set of no-signaling correlations. For example, in the $(2,2,2)$ scenario, after imposing the no-signaling and normalization constraints, the maximal dimension is reduced to $\dim P-1=8$. If we set $p(01\mid 00)$ as a zero constraint, one parameter is fixed to zero while the others remain free, resulting in a lower-dimensional face. Adding another zero constraint, such as $p(10\mid 00)=0$, further reduces the dimensionality of the face. Thus, for the input $(X_{1},X_{2})=(0,0)$, we obtain
$p(00\mid 00)+p(11\mid 00)=1$ (47)
which characterizes a lower-dimensional proper face:
$F(X_{1}=0,X_{2}=0)=\{\mathbf{p}\mid p(01\mid 00)+p(10\mid 00)=0\}$ (48)
Applying the same procedure for other inputs yields correlations such that, for given inputs $X_{1},X_{2}$ and outputs $x_{1},x_{2}$, the logical formula $x_{1}\oplus x_{2}=f(X_{1},X_{2})$ must hold with certainty, i.e., $p(x_{1}\oplus x_{2}=f(X_{1},X_{2})\mid X_{1},X_{2})=1$. Any probabilities violating this logical formula are forced to zero by the imposed constraints. Hence, the zero constraints directly correspond to the logical formula and lead to an AVN (all-versus-nothing) proof. These zeros define the face of the polytope with maximally reduced dimension.
Appendix B Parametric table for $(3,2,2)$-scenario
The reduced parametric table for the $(3,2,2)$ scenario is obtained using no-signaling and the normalization condition mentioned in Section(II).
Con.& Sec. (0,0,0) (0,0,1) (0,1,0) (0,1,1) (1,0,0) (1,0,1) (1,1,0) (1,1,1)
$(0,0,0)$ $p_{1}$ $p_{2}$ $p_{3}$ $p_{4}$ $p_{5}$ $p_{6}$ $p_{7}$ $1-(p_{1}+p_{2}+p_{3}+p_{4}+p_{5}+p_{6}+p_{7})$
$(0,0,1)$ $p_{9}$ $p_{1}+p_{2}-p_{9}$ $p_{11}$ $p_{3}+p_{4}-p_{11}$ $p_{13}$ $p_{5}+p_{6}-p_{13}$ $p_{15}$ $1-(p_{1}+p_{2}+p_{3}+p_{4}+p_{5}+p_{6}+p_{15})$
$(0,1,0)$ $p_{17}$ $p_{18}$ $p_{1}+p_{3}-p_{17}$ $p_{2}+p_{4}-p_{18}$ $p_{21}$ $p_{22}$ $p_{5}+p_{7}-p_{21}$ $1-(p_{1}+p_{2}+p_{3}+p_{4}+p_{5}+p_{7}+p_{22})$
$(0,1,1)$ $p_{25}$ $p_{17}+p_{18}-p_{25}$ $p_{9}+p_{11}-p_{25}$ $p_{1}+p_{2}+p_{3}+p_{4}-p_{9}-p_{11}-p_{17}-p_{18}+p_{25}$ $p_{8}$ $p_{21}+p_{22}-p_{8}$ $p_{13}+p_{15}-p_{8}$ $1-(p_{1}+p_{2}+p_{3}+p_{4}+p_{21}+p_{22}+p_{13}+p_{15}-p_{8})$
$(1,0,0)$ $p_{10}$ $p_{12}$ $p_{14}$ $p_{16}$ $p_{1}+p_{5}-p_{10}$ $p_{2}+p_{6}-p_{12}$ $p_{3}+p_{7}-p_{14}$ $1-(p_{1}+p_{2}+p_{3}+p_{5}+p_{6}+p_{7}+p_{16})$
$(1,0,1)$ $p_{19}$ $p_{10}+p_{12}-p_{19}$ $p_{20}$ $p_{14}+p_{16}-p_{20}$ $p_{9}+p_{13}-p_{19}$ $p_{1}+p_{2}+p_{5}+p_{6}-p_{9}-p_{13}-p_{10}-p_{12}+p_{19}$ $p_{11}+p_{15}-p_{20}$ $1-(p_{1}+p_{2}+p_{5}+p_{6}+p_{11}+p_{14}+p_{16}-p_{20})$
$(1,1,0)$ $p_{23}$ $p_{24}$ $p_{10}+p_{14}-p_{23}$ $p_{12}+p_{16}-p_{24}$ $p_{17}+p_{21}-p_{23}$ $p_{18}+p_{22}-p_{24}$ $p_{1}+p_{3}+p_{5}+p_{7}-p_{17}-p_{21}-p_{10}-p_{12}+p_{23}$ $1-(p_{1}+p_{3}+p_{5}+p_{7}+p_{18}+p_{22}+p_{12}+p_{16}-p_{24})$
$(1,1,1)$ $p_{26}$ $p_{23}+p_{24}-p_{26}$ $p_{19}+p_{20}-p_{26}$ $p_{10}+p_{12}+p_{14}+p_{16}-p_{19}-p_{20}-p_{23}-p_{24}+p_{26}$ $p_{25}+p_{8}-p_{26}$ $p_{17}+p_{18}+p_{21}+p_{22}-p_{25}-p_{8}-p_{23}-p_{24}+p_{26}$ $p_{9}+p_{11}+p_{13}+p_{15}-p_{25}-p_{8}-p_{19}-p_{20}+p_{26}$ $1-(p_{9}+p_{11}+p_{13}+p_{15}+p_{17}+p_{18}+p_{21}+p_{22}-p_{25}-p_{8}+p_{10}+p_{12}+p_{14}+p_{16}-p_{19}-p_{20}-p_{23}-p_{24}+p_{26})$
Table 4: Reduced (26 free parameters) probability table.
Appendix C Boolean no-signaling condition
The Boolean no-signaling condition is the possibilistic form of Eq.(1), such that the probability parameters are replaced by Boolean formulas ($B_{i}^{\prime}s$) and the summation operation is replaced with the Boolean OR operation.
$\bigvee_{x_{j}}B(x_{1},\cdots,x_{j},\cdots,x_{n}\mid X_{1},\cdots,X_{j},\cdots,X_{n})=\bigvee_{x_{j}}B(x_{1},\cdots,x_{j},\cdots,x_{n}\mid X_{1},\cdots,X_{j}^{\prime},\cdots,X_{n})$ (49)
for all $j$, all outcomes $x_{1},\dots,x_{j-1},x_{j+1},\dots,x_{n}$, and all settings $X_{1},\dots,X_{j-1},X_{j},X_{j}^{\prime},X_{j+1},\dots,X_{n}$.